Using the information provided, find the parallel and the perpendicular equation. Graph all three equations on a grid.
1. Y=4x+8 (2,5)
Parallel equation:
Perpendicular equation
2. y=3x-5 (6,9)
Parallel equation:

Perpendicular equation:

3. y=-5x+3 (-2,-3)
parallel equation:
Perpendicular equation:

Answer by algebrapro18(249)

To do these you need to know that parallel lines have the same slope and that perpendicular lines have opposite reciprocal slopes.

1. Y=4x+8 (2,5)
Parallel equation: y = 4x-3
well since the slope of the original line was 4 the parallel line will also have a slope of 4. Now we can find the equation of this line two ways.

Method 1: using y=mx+b and solving for b.
we know that y = 5 and m = 4 and x =2 from the given information so now we just need to calculate b.
y=mx+b
5 = 4*2 + b
5 = 8 + b
-3 = b

now we just back into the y=mx+b form and the equation of the line becomes:
y = 4x-3

Method 2: Using point-slop formula
we know that y1 = 5 and m = 4 and x1 = 2 so now we can put this into the point slope formula and solve for y.

y-y1=m(x-x1)
y - 5 = 4(x-2)
y -5 = 4x-8
y= 4x-3

either method gives the same answer.

Perpendicular equation: y = -1/4x+11/2
well since the slope of the original line was 4 the perpendicular line will have a slope of -1/4. Now we can find the equation of this line two ways.

Method 1: using y=mx+b and solving for b.
we know that y = 5 and m = -1/4 and x =2 from the given information so now we just need to calculate b.
y=mx+b
5 = (-1/4)*2 + b
5 = -1/2 + b
11/2 = b

now we just back into the y=mx+b form and the equation of the line becomes:
y = -1/4x+11/2

Method 2: Using point-slop formula
we know that y1 = 5 and m = -1/4 and x1 = 2 so now we can put this into the point slope formula and solve for y.

y-y1=m(x-x1)
y - 5 = -1/4(x-2)
y -5 = -1/4x+1/2
y= -1/4x+11/2

either method gives the same answer.


2. y=3x-5 (6,9)
Parallel equation: y = 3x-9

well since the slope of the original line was 3 the parallel line will also have a slope of 3. Now we can find the equation of this line two ways.

Method 1: using y=mx+b and solving for b.
we know that y = 9 and m = 3 and x =6 from the given information so now we just need to calculate b.
y=mx+b
9 = 3*6 + b
9 = 18 + b
-9 = b

now we just back into the y=mx+b form and the equation of the line becomes:
y = 3x-9

Method 2: Using point-slop formula
we know that y1 = 9 and m = 3 and x1 = 6 so now we can put this into the point slope formula and solve for y.

y-y1=m(x-x1)
y - 9 = 3(x-6)
y -9 = 3x-18
y= 3x-9

either method gives the same answer.

Perpendicular equation: y = -1/3x+11
well since the slope of the original line was 3 the perpendicular line will have a slope of -1/3. Now we can find the equation of this line two ways.

Method 1: using y=mx+b and solving for b.
we know that y = 9 and m = -1/3 and x = 6 from the given information so now we just need to calculate b.
y=mx+b
9 = (-1/3)*6 + b
9 = -2 + b
11 = b

now we just back into the y=mx+b form and the equation of the line becomes:
y = -1/3x+11

Method 2: Using point-slop formula
we know that y1 = 9 and m = -1/3 and x1 = 6 so now we can put this into the point slope formula and solve for y.

y-y1=m(x-x1)
y - 9 = -1/3(x-6)
y -9 = -1/3x + 2
y= -1/3x + 11

either method gives the same answer.

3. y=-5x+3 (-2,-3)

parallel equation: y = -5x-13

well since the slope of the original line was -5 the parallel line will also have a slope of -5. Now we can find the equation of this line two ways.

Method 1: using y=mx+b and solving for b.
we know that y = -3 and m = -5 and x = -2 from the given information so now we just need to calculate b.
y=mx+b
-3 = -5*-2 + b
-3 = 10 + b
-13 = b

now we just back into the y=mx+b form and the equation of the line becomes:
y = -5x-13

Method 2: Using point-slop formula
we know that y1 = -3 and m = -5 and x1 = -2 so now we can put this into the point slope formula and solve for y.

y-y1=m(x-x1)
y + 3 = -5(x+2)
y + 3 = -5x-10
y= -5x-13

either method gives the same answer.


Perpendicular equation: y = 1/5x-13/5

well since the slope of the original line was -5 the perpendicular line will have a slope of 1/5. Now we can find the equation of this line two ways.

Method 1: using y=mx+b and solving for b.
we know that y = -3 and m = 1/5 and x = -2 from the given information so now we just need to calculate b.
y=mx+b
-3 = (1/5)*-2 + b
-3 = -2/5 + b
-13/5 = b

now we just back into the y=mx+b form and the equation of the line becomes:
y = 1/5x-13/5

Method 2: Using point-slop formula
we know that y1 = -3 and m = 1/5 and x1 = -2 so now we can put this into the point slope formula and solve for y.

y-y1=m(x-x1)
y + 3 = 1/5(x+2)
y +3 = 1/5x + 2/5
y= 1/5x -13/5

either method gives the same answer.

I will leave the graphing to you.